OpenStudy (anonymous):

The half life of radon is 3.80 days what would be its decay constant?

OpenStudy (shayaan_mustafa):

Hi! Hope you will be fine :D Half life formula is, \[0.5=e ^{k*t}\] where, k=decay constant t=time substitute time, we get \[0.5=e ^{k*3.80}\] taking natural logarithm on both sides, we get, \[\ln(0.5)=k*3.80\] or, \[\ln(0.5)/3.80=k\] \[k=0.182\](ANSWER) Regards, Electronics engineer. *but I ain't good in nuclear physics :D